# From Landau two-fluid model to de Sitter Universe

### v10

@@ IX.8 Acoustic Planck constants
Returning to the two-fluid hydrodynamics, let us consider quasiparticles in Bose superfluid 4He. The flow of the liquid provides the acoustic gravity for ”relativistic” quasiparticles – phonons. The effective acoustic metric and thus the acoustic Planck constants, $\hbar_{\rm ac}$ and $h_{\rm ac}$, can be expressed in terms of the parameters of this liquid.
@@
The quantum mechanics of phonons demonstrates that it can be described by the effective acoustic Planck constants, $\hbar_{\rm ac}$ and $h_{\rm ac}$:
$\hbar_{\rm ac}^{2}=g^{00}=\frac{m_{4}}{\hbar ns}\,\,,\,\,/\!\!\!h_{\rm ac}^{2}=\hbar_{\rm ac}^{2}s^{2}=\frac{m_{4}s}{\hbar n}\,.$ (98)
The ground state of superfluid 4He represents the strongly interacting and strongly correlated ”quantum vacuum”. Two Planck mass scales $m_{4}s^{2}$ and $\hbar s/a$ are of the same order. That is why in this analogue of quantum vacuum the acoustic Planck constant $h_{\rm ac}\sim a$, i.e. the length of the effective Planck constant is on the order of the effective Planck length (the interatomic distance).
@@ IX.9 Phase boundary between Universes with different $\hbar$
The so-called multiple point principleNielsen1994 suggests that the quantum vacua are degenerate, and thus there is the boundary of the first order phase transition between our Universe and the Universe with very different properties.Nielsen2023 If so, it is not excluded that the two neighbouring Universes may have different Minkowski metric and thus different values of Planck parameters $\hbar$ and /h/\!\!\!h. If there is the thermal contact between these Universes, then in the thermal equilibrium one would have the following condition for temperatures and Planck constants:
@@
Another interesting question is what happens when $e_{\mu}^{a}\rightarrow 0$, i.e. we approach the critical point at which the interval $ds$ becomes identically zero and thus the spacetime disappears. According to Eq.(83, in this limit the both Planck constants approach infinity, $\hbar\rightarrow\infty$ and $h\rightarrow\infty$. This demonstrates the quantum correlations between the points separated by arbitrary large distances. However, there may be different limiting cases corresponding to different relations between $\hbar$, /h/\!\!\!h and $G$ in the Bronstein $cGh$ cube,Bronstein1933 ; Duff2002 see Ref. Ecker2025 and references therein.
@@ X Discussion
One possibility is to introduce the proper vacuum variable, which describes the quantum vacuum and its vacuum energy density. The working example of such variable is the Hawking 4-form field.Hawking1984 This variable was used in particular in the so-called $q$-theory.KlinkhamerVolovik2008 ; KlinkhamerVolovik2008c However, the suggested form of the $q$-theory gives rise to the cold dark matter instead of the stiff matter.KlinkhamerVolovik2017 The analogue of the cold matter is produced by the fast oscillations of the $q$-field. This $q$-theory mechanism also gives the required power-law decay of the dark energy similar to that in Eq.(37). However, this decay to the Minkowski vacuum takes place only in case of the proper choice of the value of corresponding analog of the chemical potential. This represents some form of fine tuning, while the two fluid dynamics does not require the fine tuning.
@@
On the other hand the entropy of the de Sitter state is extensive, i.e. for arbitrary volume $V$ of the de Sitter state, the entropy is proportional to the volume, $S(V)=sV$, where $s$ is the entropy density. But the integration of the entropy density over the Hubble volume demonstrates the holographic bulk-surface connection between the entropy of the Hubble volume and the Gibbons-Hawking surface entropy of the cosmological horizon, $S(V_{H})=sV_{H}=A/4G$.

### v11

@@ IX.8 Acoustic Planck constants
Returning to the two-fluid hydrodynamics, let us consider quasiparticles in Bose superfluid 4He. The flow of the liquid provides the acoustic gravity for ”relativistic” quasiparticles – phonons. The effective acoustic metric and thus the acoustic Planck constants, $\hbar_{\rm ac}$ and $\penalty 50\!\!\!h_{\rm ac}$, can be expressed in terms of the parameters of this liquid.
@@
The quantum mechanics of phonons demonstrates that it can be described by the effective acoustic Planck constants, $\hbar_{\rm ac}$ and $\penalty 50\!\!\!h_{\rm ac}$:
$\hbar_{\rm ac}^{2}=g^{00}=\frac{m_{4}}{\hbar ns}\,\,,\,\,/\penalty 50\!\!\!h_{\rm ac}^{2}=\hbar_{\rm ac}^{2}s^{2}=\frac{m_{4}s}{\hbar n}\,.$ (98)
The ground state of superfluid 4He represents the strongly interacting and strongly correlated ”quantum vacuum”. Two Planck mass scales $m_{4}s^{2}$ and $\hbar s/a$ are of the same order. That is why in this analogue of quantum vacuum the acoustic Planck constant $\penalty 50\!\!\!h_{\rm ac}\sim a$, i.e. the length of the effective Planck constant is on the order of the effective Planck length (the interatomic distance).
@@ IX.9 Phase boundary between Universes with different $\hbar$
The so-called multiple point principleNielsen1994 suggests that the quantum vacua are degenerate, and thus there is the boundary of the first order phase transition between our Universe and the Universe with very different properties.Nielsen2023 If so, it is not excluded that the two neighbouring Universes may have different Minkowski metric and thus different values of Planck parameters $\hbar$ and $\penalty 50\!\!\!h$. If there is the thermal contact between these Universes, then in the thermal equilibrium one would have the following condition for temperatures and Planck constants:
@@
Another interesting question is what happens when $e_{\mu}^{a}\rightarrow 0$, i.e. we approach the critical point at which the interval $ds$ becomes identically zero and thus the spacetime disappears. According to Eq.(83, in this limit the both Planck constants approach infinity, $\hbar\rightarrow\infty$ and $\penalty 50\!\!\!h\rightarrow\infty$. This demonstrates the quantum correlations between the points separated by arbitrary large distances. However, there may be different limiting cases corresponding to different relations between $\hbar$, $\penalty 50\!\!\!h$ and $G$ in the Bronstein $cGh$ cube,Bronstein1933 ; Duff2002 see Ref. Ecker2025 and references therein.
@@ X Discussion
One possibility is to introduce the proper vacuum variable, which describes the quantum vacuum and its vacuum energy density. The working example of such variable is the Hawking 4-form field.Hawking1984 This variable was used in particular in the so-called $q$-theory.KlinkhamerVolovik2008c ; KlinkhamerVolovik2008G However, the suggested form of the $q$-theory gives rise to the cold dark matter instead of the stiff matter.KlinkhamerVolovik2017 The analogue of the cold matter is produced by the fast oscillations of the $q$-field. This $q$-theory mechanism also gives the required power-law decay of the dark energy similar to that in Eq.(37). However, this decay to the Minkowski vacuum takes place only in case of the proper choice of the value of corresponding analog of the chemical potential. This represents some form of fine tuning, while the two fluid dynamics does not require the fine tuning.
@@
On the other hand the entropy of the de Sitter state is extensive, i.e. for arbitrary volume $V$ in the de Sitter Universe, the entropy is proportional to the volume, $S(V)=sV$, where $s$ is the entropy density. But the integration of the entropy density over the Hubble volume demonstrates the holographic bulk-surface connection between the entropy of the Hubble volume and the Gibbons-Hawking surface entropy of the cosmological horizon, $S(V_{H})=sV_{H}=A/4G$.